# Uncovering Hidden Connections: Iterative Search and Reasoning for Video-grounded Dialog

### v3

@@ 3.2.3 Path Aggregation
3.3 Visual Encoder
3.3.1 Preprocessing
Following [21, 19], we adopt Faster R-CNN model [^6] https://github.com/peteanderson80/bottom-up-attention. pretrained on the Visual Genome dataset to extract object features of each frame in the video, which is a renowned convolutional network model optimized for object detection task [44, 45, 46]. Given the richness of videos in terms of the number of frames ($F$) and the plethora of objects ($O$) within these frames, we incorporate them into a single dimension, denoted as $\mathbf{E}^{{}^{\prime}}_{v}\in\mathbb{R}^{{n_{v}}\times{2048}}$, where $n_{v}=F\times O$.
@@ 3.3.1 Preprocessing
$\mathbf{E}_{v}=\mathrm{Norm}(\phi(f_{v}(\mathbf{E}^{{}^{\prime}}_{v}))),$ (6)
@@
3.3.2 Iterative Reasoning
For a comprehensive extraction of data related to the question $Q_{t}$ from video $V$, we introduce a novel multimodal iterative reasoning network. This design prioritizes cyclical modality interactions, incrementally refining the information acquired. Both modalities converge into a weight matrix, described as:
$\begin{cases}\mathbf{E}=\mathbf{E}_{q}||\mathbf{E}_{v},\\ \mathbf{W}=\mathrm{Softmax}(\mathbf{E}\mathbf{E}^{T}),\end{cases}$ (7)
@@ 3.3.2 Iterative Reasoning
We utilize four separate attention networks, each capturing distinct semantic information and enhancing the representation across different dimensions. For example, the video-augmented question representations are derived as:
$\begin{cases}\mathbf{E}_{q}^{v}=\mathbf{W}^{O}||_{j=1}^{J}\mathrm{head}_{j},\\ \mathrm{head}_{j}=\mathrm{Attention}(\mathbf{E}_{q}\mathbf{W}_{j}^{Q},\mathbf{E}_{v}\mathbf{W}_{j}^{K},\mathbf{E}_{v}\mathbf{W}_{j}^{V}),\\ \mathrm{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{Softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}})\mathbf{V},\end{cases}$ (8)
@@
$\begin{cases}\mathbf{E}_{q}^{vis}=\mathrm{Sum}(\mathbf{W}[:n_{q},:n_{q}])*\mathbf{E}_{q}^{q}+\mathrm{Sum}(\mathbf{W}[:n_{q},n_{q}:])*\mathbf{E}_{q}^{v},\\ \hat{\mathbf{E}}_{v}=\mathrm{Sum}(\mathbf{W}[n_{q}:,:n_{q}])*\mathbf{E}_{v}^{q}+\mathrm{Sum}(\mathbf{W}[n_{q}:,n_{q}:])*\mathbf{E}_{v}^{v},\end{cases}$ (9)
@@
TABLE I: Performance comparison with the latest methods on the AVSD@DSTC7 dataset. The best results are highlighted in bold and the second is underlined.
Method Year BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
Baseline [15] 2019 0.626 0.485 0.383 0.309 0.215 0.487 0.746
HMA [47] 2019 0.633 0.490 0.386 0.310 0.242 0.515 0.856
RMFF [48] 2019 0.636 0.510 0.417 0.345 0.224 0.505 0.877
EE-DMN [49] 2019 0.641 0.493 0.388 0.310 0.241 0.527 0.912
HAN [50] 2019 0.718 0.584 0.478 0.394 0.267 0.563 1.094
MTN [16] 2019 0.731 0.597 0.490 0.406 0.271 0.564 1.127
JMAN [51] 2020 0.667 0.521 0.413 0.334 0.239 0.533 0.941
MSTN [52] 2020 - - - 0.377 0.275 0.566 1.115
MTN-P [26] 2020 0.750 0.619 0.514 0.427 0.280 0.580 1.189
BiST [23] 2020 0.755 0.619 0.510 0.429 0.284 0.581 1.192
VGNMN [21] 2021 - - - 0.429 0.278 0.578 1.188
SCGA [19] 2021 0.745 0.622 0.517 0.430 0.285 0.578 1.201
VGD [27] 2020 0.750 0.621 0.516 0.433 0.283 0.581 1.196
PDC [20] 2021 0.770 0.653 0.539 0.449 0.292 0.606 1.295
RLM [17] 2021 0.765 0.643 0.543 0.459 0.294 0.606 1.308
T5RLM [28] 2022 0.767 0.644 0.542 0.461 0.296 0.608 1.311
JVIT [24] 2022 0.776 0.652 0.544 0.453 0.300 0.614 1.315
DialogMCF [25] 2023 0.777 0.653 0.547 0.457 0.306 0.613 1.352
ISR - 0.782 0.655 0.552 0.469 0.305 0.619 1.331
3.4 Generator
@@
$\mathbf{g}=\sigma(\mathbf{W}_{g}[\mathbf{E}_{q}||\mathbf{E}_{q}^{txt}||\mathbf{E}_{q}^{vis}]),$ (10)
@@
$\hat{\mathbf{E}}_{q}=\mathbf{g}\odot\mathbf{E}_{q}^{txt}+(\mathbf{1}-\mathbf{g})\odot\mathbf{E}_{q}^{vis},$ (11)
@@
Building upon the methodology proposed by [20], we utilize the pre-trained GPT-2 model, recognized for its advanced text generation capabilities, as the backbone for our response generator. Specifically, the generator creates responses based on the representations of the video $\hat{\mathbf{E}}_{v}$, the caption $\mathbf{E}_{c}$, the dialog history $\hat{\mathbf{E}}_{h}$, and the posed question $\hat{\mathbf{E}}_{q}$. The combined input for the generator is structured as $(\hat{\mathbf{E}}_{v},[\mathrm{SEP}],\mathbf{E}_{c},[\mathrm{SEP}],\hat{\mathbf{E}}_{h},[\mathrm{SEP}],\hat{\mathbf{E}}_{q})$. Utilizing several transformer layers, the probability of generating a response $A_{t}=(w_{1}^{a},w_{2}^{a},...,w_{n_{a}}^{a})$ is derived as:
${\begin{split}p(A_{t}|V,C,H,Q_{t};\theta)&\approx p(A_{t}|\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta)\\ &=\prod_{i=1}^{n_{a}}p(w_{i}^{a}|w_{<i}^{a},\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta),\end{split}}$ (12)
@@
3.5 Optimization
To refine our optimization strategy, we focus on reducing the negative log-likelihood associated with the actual answer. Therefore, we employ the Maximum Likelihood Estimation (MLE) [17] loss and adopt the gradient descent method for optimization. The expression for the loss is:
${\mathcal{L}_{mle}=-\sum_{i=1}^{n_{a}}\mathrm{log}\,p(w_{i}^{a}|w_{<i}^{a},\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta).}$ (13)
TABLE II: Details of the benchmark datasets AVSD@DSTC7, AVSD@DSTC8, and VSTAR.
@@ 3.5 Optimization
4 Experiments
4.1 Dataset
We employed three publicly available datasets to evaluate our model, including AVSD@DSTC7 [38], AVSD@DSTC8 [39], and VSTAR [53]. The first two datasets are both derived from the Charades human-activity dataset [54] that captures authentic human interactions about video clips. The training and validation sets in both datasets are the same, while the testing set is completely different. Due to longer sentences in DSTC8, more nuanced utterance comprehension as well as long-term history dependency capture is more critical. Each dialog, tied to a video, contains ten paired question-answer utterances. VSTAR is a large-scale video-grounded dialog understanding dataset, which is collected from 395 TV series (185K 90-second clips) with carefully cleaned dialogs and metadata information. The sheer volume of data exacerbates the complexity of this dataset. Dataset specifics are detailed in Table II.
@@ 4.1 Dataset
TABLE III: Performance comparison with the latest methods on the AVSD@DSTC8 dataset. The best results are highlighted in bold and the second is underlined.
Method Year BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
Baseline [15] 2019 - - - 0.293 0.212 0.483 0.679
MTN [16] 2019 - - - 0.352 0.263 0.547 0.978
MDMN [55] 2020 - - - 0.296 0.214 0.496 0.761
JMAN [51] 2020 0.645 0.504 0.402 0.324 0.232 0.521 0.875
MSTN [52] 2020 - - - 0.385 0.270 0.564 1.073
MTN-P [26] 2020 0.701 0.587 0.494 0.419 0.263 0.564 1.097
STSGR [22] 2021 - - - 0.357 0.267 0.553 1.004
SCGA [19] 2021 0.711 0.593 0.497 0.416 0.276 0.566 1.123
RLM [17] 2021 0.746 0.626 0.528 0.445 0.286 0.598 1.240
T5RLM [28] 2022 0.749 0.631 0.529 0.445 0.290 0.600 1.263
JVIT [24] 2022 0.748 0.632 0.536 0.456 0.289 0.600 1.268
DialogMCF [25] 2023 0.756 0.633 0.532 0.449 0.293 0.601 1.253
ISR - 0.762 0.641 0.543 0.460 0.298 0.607 1.285
@@
TABLE IV: Performance comparison with the latest methods on the VSTAR dataset. The best results are highlighted in bold and the second is underlined.
@@
OpenViDial coarse [56] 0.006 0.035 0.063 0.066
RLM [17] 0.010 0.032 0.061 0.079
AVDT [53] 0.014 0.040 0.080 0.137
ISR 0.020 0.044 0.081 0.150
4.2 Experimental Settings
4.2.1 Implementation Details
For model architecture, we fine-tuned the GPT-2 generator and froze the Faster R-CNN model as well as skipping the pre-trained text embedding model. We trained our model for 20 epochs using the Adam optimizer. The learning rate was set at 4.5e-4 and reduced tenfold between the 5-th and 10-th epochs. Configurations included a frame count ($F$) of 15, object number ($O$) of 6, and attention heads ($J$) of 3. The feature dimension ($d$) was 768, and batches were organized in sets of 32. During decoding, we employed a beam search with a width of 5, a sequence length capped at 20, and a length penalty of 0.2. Hyperparameters, the selection threshold ($p$) and iteration number ($I$), were separately set to 0.6 and 3 using a grid search on the validation set, where the range of $p$ is {0.0, 0.2, 0.4, 0.6, 0.8, 1.0} and $I$ ranges from 0 to 5. Experiments were conducted using PyTorch on a single NVIDIA A100-PCIE-40GB GPU with CUDA 11.0.
4.2.2 Evaluation Metrics
4.3 Performance Comparison
We evaluated the performance of our proposed framework against several state-of-the-art baselines. This comparison includes generalized methods (e.g., MTN [16], BiST [23], SCGA [19]), as well as GPT-based ones (e.g., PDC [20], RLM [17], T5RLM [28]) on the AVSD@DSTC7 and AVSD@DSTC8 datasets. The results for the AVSD@DSTC7 dataset are summarized in Table I, with the following notable insights:
@@ 4.3 Performance Comparison
  • Our model surpassed all compared methods, achieving the highest scores across almost all evaluation metrics. For example, ISR achieves a 3.5% relative gain in BLEU-4 score compared to the JVIT baseline. These results not only validate the effectiveness of our overall framework but also reveal the necessity of mining deeper clues for question understanding.
The performance of our proposed framework, juxtaposed against several state-of-the-art methods on the AVSD@DSTC8 dataset, is delineated in Table III. Key observations drawn from the comparative results are as follows:
@@
  • Our proposed model demonstrates significant advancements, achieving improvements across multiple evaluation metrics on the AVSD@DSTC8 benchmark. Notably, it surpasses DialogMCF by 2.4% in BLEU-4 and by 2.5% in CIDEr scores. These findings underscore the robustness and enhanced effectiveness of our model.
The evaluation results of our ISR model on the VSTAR dataset are shown in Table IV. And we can derive the following conclusions from this:
@@
TABLE V: Performance comparison among the variants of our model on the AVSD@DSTC7 dataset. The best results are highlighted in bold.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
ISR 0.782 0.655 0.552 0.469 0.305 0.619 1.331
w/o textual encoder 0.774 0.644 0.539 0.455 0.299 0.612 1.289
w/o visual encoder 0.771 0.642 0.537 0.452 0.298 0.613 1.292
w/o gate mechanism 0.775 0.650 0.547 0.465 0.301 0.616 1.326
w/o GPT-2 decoder 0.760 0.636 0.528 0.441 0.287 0.603 1.229
TABLE VI: Performance comparison with the large video-language models. The version of ChatGPT is “gpt-3.5-turbo” and the best performance is highlighted in bold.
Methods Objective Metric ChatGPT
BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr Accuracy Score
AVSD@DSTC7
Video-LLaMA-7B 0.360 0.244 0.165 0.115 0.210 0.367 0.383 28.533 2.551
Video-LLaVA-7B 0.422 0.291 0.200 0.142 0.239 0.418 0.447 33.158 2.918
ISR 0.782 0.655 0.552 0.469 0.305 0.619 1.331 51.843 3.630
AVSD@DSTC8
Video-LLaMA-7B 0.345 0.231 0.158 0.110 0.201 0.364 0.444 28.737 2.575
Video-LLaVA-7B 0.410 0.280 0.194 0.137 0.228 0.416 0.527 33.275 2.998
ISR 0.762 0.641 0.543 0.460 0.298 0.607 1.285 51.012 3.588
4.4 Ablation Study
The goal of our ablation study is to critically evaluate the individual contribution of distinct modules within our ISR model. By strategically removing certain components, we can assess their significance in the holistic performance of our model. We designed four derivatives to accomplish this:

### v4

@@ 3.2.3. Path Aggregation
Figure 4. Illustration of our multimodal iterative reasoning network.
3.3. Visual Encoder
3.3.1. Preprocessing
Following (Le et al., 2022; Kim et al., 2021), we adopt Faster R-CNN model [^6] https://github.com/peteanderson80/bottom-up-attention. pretrained on the Visual Genome dataset to extract object features of each frame in the video, which is a renowned convolutional network model optimized for object detection task (Ren et al., 2015; Yin et al., 2022; Vecchio et al., 2020). Given the richness of videos in terms of the number of frames ($F$) and the plethora of objects ($O$) within these frames, we incorporate them into a single dimension, denoted as $\mathbf{E}^{{}^{\prime}}_{v}\in\mathbb{R}^{{n_{v}}\times{2048}}$, where $n_{v}=F\times O$.
@@ 3.3.1. Preprocessing
(6) $\mathbf{E}_{v}=\mathrm{Norm}(\phi(f_{v}(\mathbf{E}^{{}^{\prime}}_{v}))),$
@@
3.3.2. Iterative Reasoning
For a comprehensive extraction of data related to the question $Q_{t}$ from video $V$, we introduce a novel multimodal iterative reasoning network, as shown in Fig. 4. This design prioritizes cyclical modality interactions, incrementally refining the information acquired. Both modalities converge into a weight matrix, described as:
(7) $\begin{cases}\mathbf{E}=\mathbf{E}_{q}||\mathbf{E}_{v},\\ \mathbf{W}=\mathrm{Softmax}(\mathbf{E}\mathbf{E}^{T}),\end{cases}$
@@ 3.3.2. Iterative Reasoning
We utilize four separate attention networks (Zhang et al., 2023a) without sharing parameters, each capturing distinct semantic information and enhancing the representation across different dimensions. This setup serves two advantages: 1) Each attention network can focus on learning distinct representations for different modalities, preventing information loss due to excessive feature blending. 2) A single attention mechanism may struggle to process diverse modalities efficiently, leading to information bottlenecks. Independent networks ensure that each modality gets sufficient attention without interference. For example, the video-augmented question representations are derived as:
(8) $\begin{cases}\mathbf{E}_{q}^{v}=\mathbf{W}^{O}||_{j=1}^{J}\mathrm{head}_{j},\\ \mathrm{head}_{j}=\mathrm{Attention}(\mathbf{E}_{q}\mathbf{W}_{j}^{Q},\mathbf{E}_{v}\mathbf{W}_{j}^{K},\mathbf{E}_{v}\mathbf{W}_{j}^{V}),\\ \mathrm{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{Softmax}(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}})\mathbf{V},\end{cases}$
@@
(9) $\begin{cases}\mathbf{E}_{q}^{vis}=\mathrm{Sum}(\mathbf{W}[:n_{q},:n_{q}])*\mathbf{E}_{q}^{q}+\mathrm{Sum}(\mathbf{W}[:n_{q},n_{q}:])*\mathbf{E}_{q}^{v},\\ \hat{\mathbf{E}}_{v}=\mathrm{Sum}(\mathbf{W}[n_{q}:,:n_{q}])*\mathbf{E}_{v}^{q}+\mathrm{Sum}(\mathbf{W}[n_{q}:,n_{q}:])*\mathbf{E}_{v}^{v},\end{cases}$
@@
3.4. Generator
@@
(10) $\mathbf{g}=\sigma(\mathbf{W}_{g}[\mathbf{E}_{q}||\mathbf{E}_{q}^{txt}||\mathbf{E}_{q}^{vis}]),$
@@
(11) $\hat{\mathbf{E}}_{q}=\mathbf{g}\odot\mathbf{E}_{q}^{txt}+(\mathbf{1}-\mathbf{g})\odot\mathbf{E}_{q}^{vis},$
@@
Building upon the methodology proposed by (Le et al., 2021), we utilize the pre-trained GPT-2 model, recognized for its advanced text generation capabilities, as the backbone for our response generator. Specifically, the generator creates responses based on the representations of the video $\hat{\mathbf{E}}_{v}$, the caption $\mathbf{E}_{c}$, the dialog history $\hat{\mathbf{E}}_{h}$, and the posed question $\hat{\mathbf{E}}_{q}$. The combined input for the generator is structured as $(\hat{\mathbf{E}}_{v},[\mathrm{SEP}],\mathbf{E}_{c},[\mathrm{SEP}],\hat{\mathbf{E}}_{h},[\mathrm{SEP}],\hat{\mathbf{E}}_{q})$. Utilizing several transformer layers, the probability of generating a response $A_{t}=(w_{1}^{a},w_{2}^{a},...,w_{n_{a}}^{a})$ is derived as:
(12) ${\begin{split}p(A_{t}|V,C,H,Q_{t};\theta)&\approx p(A_{t}|\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta)\\ &=\prod_{i=1}^{n_{a}}p(w_{i}^{a}|w_{<i}^{a},\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta),\end{split}}$
@@
3.5. Optimization
To refine our optimization strategy, we focus on reducing the negative log-likelihood associated with the actual answer. Therefore, we employ the Maximum Likelihood Estimation (MLE) (Li et al., 2021b; Lan et al., 2024) loss and adopt the gradient descent method for optimization. The expression for the loss is:
(13) ${\mathcal{L}_{mle}=-\sum_{i=1}^{n_{a}}\mathrm{log}\,p(w_{i}^{a}|w_{<i}^{a},\hat{\mathbf{E}}_{v},\mathbf{E}_{c},\hat{\mathbf{E}}_{h},\hat{\mathbf{E}}_{q};\theta).}$
Table 1. Details of the benchmark datasets AVSD@DSTC7, AVSD@DSTC8, and VSTAR.
@@ 3.5. Optimization
Table 2. Performance comparison with the latest methods on the AVSD@DSTC7 dataset. The best results are highlighted in bold and the second is underlined.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
Baseline (Hori et al., 2019) 0.626 0.485 0.383 0.309 0.215 0.487 0.746
HMA (Le et al., 2019a) 0.633 0.490 0.386 0.310 0.242 0.515 0.856
RMFF (Yeh et al., 2019) 0.636 0.510 0.417 0.345 0.224 0.505 0.877
EE-DMN (Lin et al., 2019) 0.641 0.493 0.388 0.310 0.241 0.527 0.912
HAN (Sanabria et al., 2019) 0.718 0.584 0.478 0.394 0.267 0.563 1.094
MTN (Le et al., 2019b) 0.731 0.597 0.490 0.406 0.271 0.564 1.127
JMAN (Chu et al., 2020) 0.667 0.521 0.413 0.334 0.239 0.533 0.941
MSTN (Lee et al., 2020) - - - 0.377 0.275 0.566 1.115
MTN-P (Le and Chen, 2020) 0.750 0.619 0.514 0.427 0.280 0.580 1.189
BiST (Le et al., 2020) 0.755 0.619 0.510 0.429 0.284 0.581 1.192
VGNMN (Le et al., 2022) - - - 0.429 0.278 0.578 1.188
SCGA (Kim et al., 2021) 0.745 0.622 0.517 0.430 0.285 0.578 1.201
VGD (Le and Hoi, 2020) 0.750 0.621 0.516 0.433 0.283 0.581 1.196
PDC (Le et al., 2021) 0.770 0.653 0.539 0.449 0.292 0.606 1.295
RLM (Li et al., 2021b) 0.765 0.643 0.543 0.459 0.294 0.606 1.308
T5RLM (Yoon et al., 2022) 0.767 0.644 0.542 0.461 0.296 0.608 1.311
JVIT (Zhang et al., 2022) 0.776 0.652 0.544 0.453 0.300 0.614 1.315
DialogMCF (Chen et al., 2023) 0.777 0.653 0.547 0.457 0.306 0.613 1.352
ISR 0.792 0.665 0.560 0.472 0.312 0.623 1.344
4. Experiments
4.1. Dataset
We employed three publicly available datasets to evaluate our model, including AVSD@DSTC7 (Alamri et al., 2018), AVSD@DSTC8 (Hori et al., 2020), and VSTAR (Wang et al., 2023). The first two datasets are both derived from the Charades human-activity dataset (Sigurdsson et al., 2016) that captures authentic human interactions about video clips. The training and validation sets in both datasets are the same, while the testing set is completely different. Due to longer sentences in DSTC8, more nuanced utterance comprehension as well as long-term history dependency capture is more critical. Each dialog, tied to a video, contains ten paired question-answer utterances. VSTAR is a large-scale video-grounded dialog understanding dataset, which is collected from 395 TV series (185K 90-second clips) with carefully cleaned dialogs and metadata information. The sheer volume of data exacerbates the complexity of this dataset. Dataset specifics are detailed in Table 1.
4.2. Experimental Settings
4.2.1. Implementation Details
For model architecture, we fine-tuned the GPT-2 generator and froze the Faster R-CNN model as well as skipping the pre-trained text embedding model. We trained our model for 20 epochs using the Adam optimizer. The learning rate was set at 3.5e-4 and reduced tenfold between the 5-th and 10-th epochs. Configurations included a frame count ($F$) of 15, object number ($O$) of 6, and attention heads ($J$) of 3. The feature dimension ($d$) was 768, and batches were organized in sets of 32. During decoding, we employed a beam search with a width of 5, a sequence length capped at 20, and a length penalty of 0.2. Hyperparameters, the selection threshold ($p$) and iteration number ($I$), were separately set to 0.6 and 3 using a grid search on the validation set, where the range of $p$ is {0.0, 0.2, 0.4, 0.6, 0.8, 1.0} and $I$ ranges from 0 to 5. Experiments were conducted using PyTorch on a single NVIDIA A100-PCIE-40GB GPU with CUDA 11.0.
4.2.2. Evaluation Metrics
@@
Table 3. Performance comparison with the latest methods on the AVSD@DSTC8 dataset. The best results are highlighted in bold and the second is underlined.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr
Baseline (Hori et al., 2019) - - - 0.293 0.212 0.483 0.679
MTN (Le et al., 2019b) - - - 0.352 0.263 0.547 0.978
MDMN (Xie and Iacobacci, 2020) - - - 0.296 0.214 0.496 0.761
JMAN (Chu et al., 2020) 0.645 0.504 0.402 0.324 0.232 0.521 0.875
MSTN (Lee et al., 2020) - - - 0.385 0.270 0.564 1.073
MTN-P (Le and Chen, 2020) 0.701 0.587 0.494 0.419 0.263 0.564 1.097
STSGR (Geng et al., 2021) - - - 0.357 0.267 0.553 1.004
SCGA (Kim et al., 2021) 0.711 0.593 0.497 0.416 0.276 0.566 1.123
RLM (Li et al., 2021b) 0.746 0.626 0.528 0.445 0.286 0.598 1.240
T5RLM (Yoon et al., 2022) 0.749 0.631 0.529 0.445 0.290 0.600 1.263
JVIT (Zhang et al., 2022) 0.748 0.632 0.536 0.456 0.289 0.600 1.268
DialogMCF (Chen et al., 2023) 0.756 0.633 0.532 0.449 0.293 0.601 1.253
ISR 0.767 0.645 0.548 0.464 0.304 0.610 1.288
Table 4. Performance comparison with the latest methods on the VSTAR dataset. The best results are highlighted in bold and the second is underlined.
@@
OpenViDial coarse (Wang et al., 2021) 0.006 0.035 0.063 0.066
RLM (Li et al., 2021b) 0.010 0.032 0.061 0.079
AVDT (Wang et al., 2023) 0.014 0.040 0.080 0.137
ISR 0.026 0.051 0.098 0.157
4.3. Performance Comparison
We evaluated the performance of our proposed framework against several state-of-the-art baselines. This comparison includes generalized methods (e.g., MTN (Le et al., 2019b), BiST (Le et al., 2020), SCGA (Kim et al., 2021)), as well as GPT-based ones (e.g., PDC (Le et al., 2021), RLM (Li et al., 2021b), T5RLM (Yoon et al., 2022)) on the AVSD@DSTC7 and AVSD@DSTC8 datasets. The results for the AVSD@DSTC7 dataset are summarized in Table 2, with the following notable insights:
@@ 4.3. Performance Comparison
  • Our model surpassed all compared methods, achieving the highest scores across almost all evaluation metrics. For example, ISR achieves a 4.2% relative gain in BLEU-4 score compared to the JVIT baseline. These results not only validate the effectiveness of our overall framework but also reveal the necessity of mining deeper clues for question understanding.
The performance of our proposed framework, in comparison with several state-of-the-art methods on the AVSD@DSTC8 dataset, is delineated in Table 3. Key observations drawn from the comparative results are as follows:
@@
  • Our proposed model demonstrates significant advancements, achieving improvements across multiple evaluation metrics on the AVSD@DSTC8 benchmark. Notably, it surpasses DialogMCF by 3.3% in BLEU-4 and by 2.8% in CIDEr scores. These findings underscore the robustness and enhanced effectiveness of our model.
The evaluation results of our ISR model on the VSTAR dataset are shown in Table 4. And we can derive the following conclusions from this:
@@
Table 5. Performance comparison among the variants of our model on the AVSD@DSTC7 dataset. The best results are highlighted in bold.
Method BLEU-1 BLEU-4 METEOR ROUGE-L CIDEr
ISR 0.792 0.472 0.315 0.623 1.344
w/o textual encoder 0.779 0.455 0.306 0.614 1.311
w/ single attention 0.785 0.467 0.311 0.619 1.334
w/o weight matrix 0.788 0.469 0.313 0.620 1.335
w/o visual encoder 0.781 0.465 0.308 0.617 1.329
w/o gate mechanism 0.786 0.468 0.310 0.620 1.338
w/o GPT-2 decoder 0.774 0.447 0.298 0.609 1.276
4.4. Ablation Study
The goal of our ablation study is to critically evaluate the individual contribution of distinct modules within our ISR model. By strategically removing certain components, we can assess their significance in the holistic performance of our model. We designed multiple derivatives to accomplish this: